Samir received $500 as a graduation gift this year and plans to deposit in an investment account. If the account pays interest of 6 percent, how much will be in his account in three years? Round your final answer to the nearest whole dollar.
Answer #1

The answer has been presented in the supporting sheet. All the parts has been solved with detailed explanation and calculation. For detailed answer refer to the supporting sheet.

В Answer 3 Future Value = amount invested *(1+i)^n 4 = 500*(1.06)^3 5 = 500*1.191016 6 = $596 8 Thus the amount in his accoun